OpenStudy (laronjaxon):

What is the equation of the line that passes through the point (2, 3) and is perpendicular to the line y = –x – 3? y = x + 5 y = –x + 5 y = x + 1 y = –x – 1

OpenStudy (anonymous):

y=x+1

OpenStudy (laronjaxon):

Lol thanks alot. so fast

OpenStudy (anonymous):

y=-x +5 and y=x+1

OpenStudy (anonymous):

@oguzcan the line y=-x+1 is not perpendicular to given line...

OpenStudy (anonymous):

sorry y=-x+5